OpenStudy (anonymous):

Fan and Medal! Helpp! Determine two different values of “b” in x2 + bx + 30 so that the expression can be factored into the product of two binomials. Explain how you determined those values and show each factorization. Explain how your process would change if the expression was 2x2 + bx + 30.

OpenStudy (anonymous):

yo

OpenStudy (anonymous):

What you want, is to open two sets of parenthesis, so that x is in the first position, and since the constant is 30, you could have (+3)(+10), then you would have (x + 3)(x+10) = x^2 + 13x + 30, so b = 13 in this case. Or, you can have (x + 6)(x + 5) which also yields x^ + bx + 30. And specifically, you would have x^2 + 11x + 30, so b = 11 in this case. Other possibilities are possible. So 2 possible values for b is b = 10 and b = 11.

OpenStudy (anonymous):

When you want to factor x^2 + bx + 30, you are looking for two numbers that when you multiply them you get +30, and when you add them up, you get b.

OpenStudy (cwrw238):

ax^2 + bx + c = 0 you needd b^2 - 4ac to be a perfect square so that the expression can be factored

OpenStudy (anonymous):

@yeah_i_suck_at_math is my answer right or wrong

OpenStudy (cwrw238):

x^2 + bx + 30 = 0 b^2 - 4ac = b^2 - 4*1*30 = b^2 - 120 we also need factors of + 30 try 5 and 6 this willl give a value of 11 and b^2- 4acc = 121 - 120 = 1 which is perfect square one solution is b = 11

OpenStudy (anonymous):

i dont know.. im the one who needs help

OpenStudy (anonymous):

okay can u try wat i said?

OpenStudy (anonymous):

no.. i didnt understand any of it, also, dont bother explaining to be honest, i still wont understand. but thanks for your attempt.

OpenStudy (anonymous):

@cwrw238 is there suppose to be another solution ...?

OpenStudy (cwrw238):

yes b = 13 is another solution as zendaya said her method of solving the first part was a good one

OpenStudy (cwrw238):

for the second part you could have (2x + 10)(x + 3) as factors which would make b = 10 + 3*2 = 16 or (2x + 3)(x + 10) giving b = 2*10 + 3 = 23

OpenStudy (cwrw238):

in the second part you have to bring 2 into the calculation of the middle term bx.